OpenStudy (aaronandyson):

A vec + B vec + C vec = 0. And A,B& C vectors are unit vectors. Find : A.B + B.C + A.C

OpenStudy (nincompoop):

here we go VECTORS

OpenStudy (aaronandyson):

A = 1 B = 1 C = 1

OpenStudy (aaronandyson):

\[1 \cos \alpha + 1 \cos \beta + 1 \cos \gamma\]